Three vectors are given by a = 2.0 i + (0) j  + (0) k  ,  b = 1.0 i + (3.0) j  + (1.0) k , and
c = ?4.4 i + (?2.0) j  + ((?4.0) k .  Find:
i)    a
Answer #1

BX C = (1.0 i + (3.0) j + (1.0) k ) X ( ?4.4 i + (?2.0) j + ((?4.0) k)

= -2 k +4j +13.2k +12i -4.4j +2i

= 14i -.4j + 11.2k

A.(BxC) = (2.0 i + (0) j + (0) k). (14i -.4j + 11.2k) = 28


A.(B+C) = (2.0 i + (0) j + (0) k).(1.0 i + (3.0) j + (1.0) k + ?4.4 i + (?2.0) j + ((?4.0) k)

= (2.0 i + (0) j + (0) k). (-3.4 i + j -3 k)

= -6.4


Ax (B+C) = (2.0 i + (0) j + (0) k) X (-3.4 i + j -3 k)

= 2.0 k + 6j


x component = 0

y component = 6

z component = 2
Answer #2

1.) ( 2i + 0j + 0k ) . ( 14 i - 0.4 k - 15.2 k )

= 28

2.) (2i ) . ( -3.4 i + j - 3 k)

= -6.8

3.) (2i) x ( -3.4 i + j - 3k)

= 6j + 2k

x component = 0

y = 6

z = 2